OpenStudy (anonymous):

True or false. Linear algebra. Similar matrices always have exactly the same eigenvalues.

jimthompson5910 (jim_thompson5910):

If A and B are similar matrices, then B = P^(-1)AP for some invertible matrix P Solve for matrix A: B = P^(-1)AP BP^(-1) = P^(-1)A PBP^(-1) = A A= PBP^(-1) Also, if k is an eigenvalue of the matrix A, then Ax = kx PBP^(-1)x = kx BP^(-1)x = P^(-1)(kx) BP^(-1)x = k*P^(-1)x which shows us that P^(-1)x is an eigenvector with eigenvalue k We can swap the roles of A and B and get a similar result. So k is an eigvenvalue for both similar matrices A and B So the statement is true.